OpenStudy (anonymous):

Indefinite integral? evaluate this indefinite integral. Integral (1+3t) t^2 dt

OpenStudy (anonymous):

what theorem am I supposed to use here?

OpenStudy (luigi0210):

You could take it apart and integrate separately? \[\LARGE \int~(1+3t)~dt~*~\int~t^2~dt\]

OpenStudy (loser66):

distribute the integrand \[\int (1+3t)(t^2) dt=\int (t^2 +3t^3)dt\] take integral term by term. Dat sit, why do you need a theorem to solve it?

OpenStudy (luigi0210):

^That works too xD

OpenStudy (luigi0210):

Sorry, haven't done real integrating in about a year ._.

OpenStudy (anonymous):

I am a little confused, how did you get t^2/3 + C , I mean what happened to the "3t^2" ??

OpenStudy (anonymous):

And I'm sorry I thought I needed a theorem thats what they were telling me but I didn't know how

OpenStudy (loser66):

Ok, let me redo it for you \[\int (1+3t)(t^2) dt=\int (t^2 +3t^3)dt\\=\frac{t^3}{3}+\frac{3t^4}{4}+C\]

OpenStudy (anonymous):

Ok now I understand. thank you!! can you help me with one more integral??

OpenStudy (loser66):

just post. I am not sure if I can help or not.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thanks anyways!!